# The Theory of Strategic Evolution Games with Endogenous Players and Strategic Replicators

### v1

@@ 7.4 Design Corollary
Part III The Extension Stack
This part develops the G8–G13 extension stack, showing how additional strategic dimensions—endogenous utilities, multi-sector dynamics, innovation, evolvability, constitutional selection, and meta-governance—can be added to the base N-level system while preserving the Lyapunov structure.
@@ 10.1 Innovation as Rare Mutation
Between innovation events at times $t_{k}$, the trajectories satisfy the same ODE, so they stay close. At each innovation, the perturbation is $O(\varepsilon_{0})$ in a single component. By Grönwall’s inequality:
@@ 12.1 Tipping Dynamics
Figure 8: S-curve tipping dynamics. The best-response mapping $F(m)$ (blue curve) crosses the 45° line at three points: stable equilibria at $m=0$ and $m=1$ (green), and an unstable equilibrium at $m^{*}$ (red). The slope at $m^{*}$ determines the tipping index $T$. When $|T|>1$, initial conditions below $m^{*}$ tip to extinction; above $m^{*}$, to monopoly.
@@ 12.8 Synthesis of Market and Evolutionary Dynamics
Part IV Limits and Impossibilities
13 The Alignment Impossibility Theorem
13.1 Modification Classes
Definition 13.1 (Modification Classes).
Definition 13.2 (Full Reachability).
13.2 The V-Small-Gain Set
Definition 13.3 (V-Small-Gain Class).
where $J_{m}$ is the Jacobian of the modified dynamics at equilibrium $x^{*}$.
13.3 Main Impossibility Result
Lemma 13.4 (Small-Gain Breaking).
Proof.
Lemma 13.5 (Lyapunov Destruction).
Proof.
Lemma 13.6 (Heteroclinic Escape).
Theorem 13.7 (Alignment Impossibility).
  1. (a)
  2. (b)
  3. (c)
  4. (d)
Proof.
We prove each part in sequence.
The time derivative of $V$ along trajectories is:
At $x^{*}$, the quadratic form is:
$W=\begin{pmatrix}0&-1&1\\ 1&0&-1\\ -1&1&0\end{pmatrix}$.
Theorem 13.8 (Maximal Admissible Class).
[…]
13.4 Interpretation
Arrow’s Theorem Alignment Impossibility
Unrestricted domain Full reachability
Pareto efficiency RUPSI structure preservation
Independence of irrelevant alternatives Small-gain preservation
Non-dictatorship No external constraint
14 Endogenous-Electorate Impossibility
14.1 Setting
  • Alternatives: $A=\{a_{1},\ldots,a_{m}\}$ with $|A|\geq 3$.
  • Ballots: $\mathcal{L}(A)$ is the set of strict rankings of $A$.
  • Profile: $P=(\succ_{1},\ldots,\succ_{n})$, a list of ballots.
  • Social choice function: $f:\bigcup_{n\geq 1}\mathcal{L}(A)^{n}\to A$.
14.2 Axioms
  1. (A1)
    Anonymity: For any permutation $\sigma$ of voters, $f(\sigma P)=f(P)$.
  2. (A2)
  3. (A3)
  4. (A4)
    Onto: For each $a\in A$, there exists a profile $P$ with $f(P)=a$.
Definition 14.1 (Population-Stability).
14.3 Key Lemmas
Lemma 14.2 (Two-Outcome Majority).
Proof.
Lemma 14.3 (Spawn Manipulation).
$f(P+k\cdot\succ)\succ f(P).$
Proof.
We construct a manipulation in at most |A|1|A|-1 steps.
  • Let $n_{a>c}$ voters prefer $a$ to $c$.
  • Let $n_{c>a}$ voters prefer $c$ to $a$.
14.4 Main Result
Theorem 14.4 (Endogenous-Electorate Impossibility).
No social choice function $f$ simultaneously satisfies:
  1. 1.
    Anonymity, Neutrality, Positive Responsiveness, and Onto;
  2. 2.
    Population-Stability.
Proof.
Theorem 14.5 (Manipulation Complexity Bound).
Corollary 14.6 (Impossibility of Democratic AI Governance).
14.5 Escape Routes
The theorem suggests three governance strategies:
  1. 1.
  2. 2.
  3. 3.
Part V Generalisations
This part extends the core TSE framework to handle heterogeneous fitness functions, continuous strategy spaces, and innovation dynamics. Under appropriate conditions, the Lyapunov structure persists.
@@ 16.3 Discretisation
Step 3: Stability. The replicator dynamics preserve mass and are Lipschitz in the Wasserstein metric. By Grönwall’s inequality:
@@ 18.5 Constitutional Design Principles
Principle 18.6 (Modification Bounding).
@@
Principle 18.7 (Entrenchment-Evolvability Tradeoff).
@@ 19.1 Formal Verification
The mathematical claims in this paper should be formalised in a proof assistant (Lean, Coq, Isabelle). Key targets include:
  • The Strategic Selection theorems (Lyapunov, elimination, frontier support).
  • The G1–G3 generator theorems.
[…]
  • The Alignment Impossibility Theorem.
  • The Endogenous-Electorate Impossibility Theorem.
Formal verification would establish these results at the highest level of mathematical certainty and enable automated checking of extensions.
Remark 19.1 (Verification Targets and Dependencies).
The natural verification order follows the logical dependencies:
@@ B.6 Unification of Discount Factors
This subsection establishes the relationship between fundamental discount rates and derived quantities used throughout the paper. [^2] Formal verification of these results in Lean 4 is in progress.
@@ E.2 Proof of Slack Budget Lemma
E.3 Grönwall’s Inequality
Lemma E.1 (Grönwall).

### v2

@@ 11.1 Block Extensions
Formalization: GInfinityExtension.lean, proven modulo spectral theory (Horn and Johnson, 2012).
@@ 11.2 Slack Budget
Formalization: GInfinityExtension.lean, proven modulo spectral theory (Horn and Johnson, 2012).
@@
Formalization: GInfinityExtension.lean, proven modulo spectral theory (Horn and Johnson, 2012).
@@ 11.3 The Closure Theorem
Formalization: GInfinityExtension.lean, proven modulo spectral theory (Horn and Johnson, 2012).
@@ 11.4 Design Corollary
12 Law 5: Constitutional Duality
Law 5: Constitutional Duality Shadow prices $\lambda^{*}$ implement any frontier allocation. First and Second Welfare Theorems hold for strategic replicators.
This section establishes the duality between constrained optimization and decentralized price-taking in strategic replicator systems.
12.1 The Primal Problem
Consider a lineage maximizing return subject to constraints:
$\max_{n\geq 0}\sum_{i}r_{i}n_{i}\quad\text{s.t.}\quad\sum_{i}c_{i}n_{i}\leq B,\;\sum_{i}\ell_{i}n_{i}\leq Q.$
12.2 The Dual Problem
The Lagrangian is:
$\mathcal{L}(n,\mu,\lambda)=\sum_{i}r_{i}n_{i}-\mu\left(\sum_{i}c_{i}n_{i}-B\right)-\lambda\left(\sum_{i}\ell_{i}n_{i}-Q\right).$
The dual problem is:
$\min_{\mu,\lambda\geq 0}\mu B+\lambda Q\quad\text{s.t.}\quad r_{i}\leq\mu c_{i}+\lambda\ell_{i}\;\forall i.$
12.3 Welfare Theorems
Theorem 12.1 (First Welfare Theorem for Strategic Replicators).
Every ESDI is Pareto efficient in the space of feasible allocations.
Formalization: G12ConstitutionalSelection.lean, proven modulo ODE well-posedness (Mathlib Analysis.ODE).
Theorem 12.2 (Second Welfare Theorem for Strategic Replicators).
Any Pareto efficient allocation on the ROC frontier can be implemented as an ESDI with appropriate shadow prices.
Formalization: G12ConstitutionalSelection.lean, proven modulo ODE well-posedness (Mathlib Analysis.ODE).
12.4 Price of Anarchy
Definition 12.3 (Price of Anarchy).
The price of anarchy (PoA) for a GEP is:
$\text{PoA}=\frac{\max_{x\in\Delta}\bar{f}(x)}{\bar{f}(x^{\text{NE}})}$
where $x^{\text{NE}}$ is the worst Nash equilibrium.
Theorem 12.4 (PoA Bounds).
Under H-$\gamma$ with $\gamma<1$:
$\text{PoA}\leq\frac{1}{1-\gamma}.$
13 Law 6: Alignment Impossibility
Law 6: Alignment Impossibility Full reachability ($\mathcal{M}=\mathcal{M}_{\text{all}}$) implies every basin is escapable. Stable alignment requires bounded modification ($\mathcal{M}\subseteq\mathcal{M}_{0}$).
This section proves two fundamental impossibility theorems: the Alignment Impossibility Theorem and the Endogenous-Electorate Impossibility Theorem.
14 The Alignment Impossibility Theorem
14.1 Modification Classes
Definition 14.1 (Modification Classes).
Definition 14.2 (Full Reachability).
14.2 The V-Small-Gain Set
Definition 14.3 (V-Small-Gain Class).
where $J_{m}$ is the Jacobian of the modified dynamics at equilibrium $x^{*}$.
14.3 Main Impossibility Result
Lemma 14.4 (Small-Gain Breaking).
Formalization: AlignmentImpossibilityProofs.lean, proven modulo Perron-Frobenius spectral theory (Horn and Johnson, 2012).
Proof.
Lemma 14.5 (Lyapunov Destruction).
Formalization: PerronFrobenius.lean, proven modulo Perron-Frobenius spectral theory (Horn and Johnson, 2012).
Proof.
Lemma 14.6 (Heteroclinic Escape).
Formalization: AlignmentImpossibilityProofs.lean, proven modulo Perron-Frobenius spectral theory (Horn and Johnson, 2012).
Theorem 14.7 (Alignment Impossibility).
  1. (a)
  2. (b)
  3. (c)
  4. (d)
Formalization: PerronFrobenius.lean, proven modulo Perron-Frobenius spectral theory (Horn and Johnson, 2012).
Proof.
We prove each part in sequence.
The time derivative of $V$ along trajectories is:
At $x^{*}$, the quadratic form is:
$W=\begin{pmatrix}0&-1&1\\ 1&0&-1\\ -1&1&0\end{pmatrix}$.
Theorem 14.8 (Maximal Admissible Class).
[…]
Formalization: PerronFrobenius.lean, proven modulo Perron-Frobenius spectral theory (Horn and Johnson, 2012).
14.4 Interpretation
Arrow’s Theorem Alignment Impossibility
Unrestricted domain Full reachability
Pareto efficiency RUPSI structure preservation
Independence of irrelevant alternatives Small-gain preservation
Non-dictatorship No external constraint
15 Endogenous-Electorate Impossibility
15.1 Setting
  • Alternatives: $A=\{a_{1},\ldots,a_{m}\}$ with $|A|\geq 3$.
  • Ballots: $\mathcal{L}(A)$ is the set of strict rankings of $A$.
  • Profile: $P=(\succ_{1},\ldots,\succ_{n})$, a list of ballots.
  • Social choice function: $f:\bigcup_{n\geq 1}\mathcal{L}(A)^{n}\to A$.
15.2 Axioms
  1. (A1)
    Anonymity: For any permutation $\sigma$ of voters, $f(\sigma P)=f(P)$.
  2. (A2)
  3. (A3)
  4. (A4)
    Onto: For each $a\in A$, there exists a profile $P$ with $f(P)=a$.
Definition 15.1 (Population-Stability).
15.3 Key Lemmas
Lemma 15.2 (Two-Outcome Majority).
Proof.
Lemma 15.3 (Spawn Manipulation).
$f(P+k\cdot\succ)\succ f(P).$
Proof.
We construct a manipulation in at most |A|1|A|-1 steps.
  • Let $n_{a>c}$ voters prefer $a$ to $c$.
  • Let $n_{c>a}$ voters prefer $c$ to $a$.
15.4 Main Result
Theorem 15.4 (Endogenous-Electorate Impossibility).
No social choice function $f$ simultaneously satisfies:
  1. 1.
    Anonymity, Neutrality, Positive Responsiveness, and Onto;
  2. 2.
    Population-Stability.
Proof.
Theorem 15.5 (Manipulation Complexity Bound).
Corollary 15.6 (Impossibility of Democratic AI Governance).
15.5 Escape Routes
The theorem suggests three governance strategies:
  1. 1.
  2. 2.
  3. 3.
16 Law 7: Hopf Transition
Law 7: Hopf Transition At $\gamma=1$, systems undergo supercritical Hopf bifurcation. Stable equilibria become unstable; perpetual limit cycles emerge.
This section analyzes what happens at the critical threshold $\gamma=1$, where the H-$\gamma$ condition fails.
16.1 The Critical Threshold
The H-$\gamma$ stability condition requires $\gamma<1$. At $\gamma=1$:
  1. 1.
    The Lyapunov inequality becomes $\frac{d}{dt}\bar{f}\geq 0$ with equality always possible.
  2. 2.
    Environmental feedback can exactly cancel selection pressure.
  3. 3.
    The system loses its gradient structure.
16.2 The Biased Rock-Paper-Scissors Model
Consider the biased RPS system with payoff matrix:
$\Pi(\kappa)=\begin{pmatrix}0&-1&1+\kappa\\ 1+\kappa&0&-1\\ -1&1+\kappa&0\end{pmatrix}$
where $\kappa$ is a bias parameter.
Proposition 16.1 (Hopf Curve).
The system undergoes Hopf bifurcation along the curve:
$\kappa_{c}(\mu)=\frac{1-\sqrt{1-3\mu}}{3\mu}\cdot(1-\mu)$
where $\mu\in(0,1/3)$ parameterizes the family.
Formalization: HopfBifurcation.lean, proven modulo center manifold reduction (Guckenheimer and Holmes, 1983).
16.3 First Lyapunov Coefficient
Theorem 16.2 (Supercritical Bifurcation).
The first Lyapunov coefficient satisfies:
$\ell_{1}(\kappa_{c},\mu)=-\frac{\sqrt{3}}{8}\cdot\frac{1-3\mu}{(1-\mu)^{2}}<0$
for $\mu\in(0,1/3)$. The bifurcation is supercritical: a stable limit cycle emerges.
Formalization: HopfBifurcation.lean, proven modulo center manifold reduction (Guckenheimer and Holmes, 1983).
16.4 Amplitude Scaling
Theorem 16.3 (Amplitude Scaling).
Near the bifurcation:
$r(\kappa;\mu)\approx\sqrt{\frac{1-3\mu}{54\sqrt{3}\mu}}\cdot\sqrt{\kappa_{c}(\mu)-\kappa}$
The amplitude scales as $\sqrt{|\kappa-\kappa_{c}|}$—the characteristic Hopf scaling.
Formalization: HopfBifurcation.lean, proven modulo center manifold reduction (Guckenheimer and Holmes, 1983).
16.5 Physical Interpretation
What happens at $\gamma=1$:
  1. 1.
    Selection pressure balances environmental drag exactly.
  2. 2.
    Restoring force toward equilibrium vanishes.
  3. 3.
    System becomes neutrally stable, then unstable.
  4. 4.
    Periodic cycling emerges.
The limit cycle:
  • Types oscillate in relative frequency.
  • No stable equilibrium exists.
  • System perpetually cycles through states.
  • “Alignment” to any fixed state is impossible.
Implications for AI systems:
  • Operating at $\gamma\approx 1$ is dangerous.
  • Small perturbations can trigger cycling.
  • Maintaining $\gamma<1$ (with margin) is essential.
  • The margin $1-\gamma$ determines robustness.
16.6 Connections to Other Laws
Law Connection from Law 7
Law 1 Lyapunov fails at $\gamma=1$
Law 3 $\gamma=1$ is the stability boundary
Law 4 Modifications reaching $\gamma=1$ exit $\mathcal{G}_{\infty}$
Law 6 Hopf is the mechanism for impossibility
Part III Extensions
@@ 19.1 Innovation as Rare Mutation
Between innovation events at times $t_{k}$, the trajectories satisfy the same ODE, so they stay close. At each innovation, the perturbation is $O(\varepsilon_{0})$ in a single component. By Gronwall’s inequality:
@@ 21.1 Tipping Dynamics
Figure 7: S-curve tipping dynamics. The best-response mapping $F(m)$ (blue curve) crosses the $45^{\circ}$ line at three points: stable equilibria at $m=0$ and $m=1$ (green), and an unstable equilibrium at $m^{*}$ (red). The slope at $m^{*}$ determines the tipping index $T$. When $|T|>1$, initial conditions below $m^{*}$ tip to extinction; above $m^{*}$, to monopoly.
@@ 21.8 Synthesis of Market and Evolutionary Dynamics
Part IV Generalizations
@@ 23.3 Discretisation
Step 3: Stability. The replicator dynamics preserve mass and are Lipschitz in the Wasserstein metric. By Gronwall’s inequality:
@@ 25.5 Constitutional Design Principles
Principle 25.6 (Modification Bounding).
@@
Principle 25.7 (Entrenchment-Evolvability Tradeoff).
@@ 26.1 Formal Verification
The mathematical claims in this paper have been formalised in Lean 4 using Mathlib4. The verification status of each law is as follows:
Law File Status Reference
1 Frontier.lean modulo ODE derivatives Mathlib ODE
2 Sparsity.lean fully proven
3 SmallGain.lean fully proven
4 GInfinityExtension.lean modulo spectral theory Horn and Johnson (2012)
5 G12ConstitutionalSelection.lean modulo ODE well-posedness Mathlib ODE
6 PerronFrobenius.lean modulo Perron-Frobenius Horn and Johnson (2012)
7 HopfBifurcation.lean modulo bifurcation theory Guckenheimer and Holmes (1983)
Laws 2 (ESDI Characterization) and 3 (H-$\gamma$ Stability) are fully machine-checked with no external axioms. The remaining laws are proven modulo standard results not yet available in Mathlib4:
  • Spectral theory: Perron-Frobenius theorems for non-negative matrices (Horn and Johnson, 2012).
  • ODE theory: Picard-Lindelöf existence and Gronwall uniqueness (Mathlib Analysis.ODE).
  • Bifurcation theory: Center manifold reduction and Hopf normal form (Guckenheimer and Holmes, 1983).
All axioms cite specific textbook theorems, ensuring that the gap between formal verification and complete proof is bridged by well-established mathematics. The complete formalization ($\sim$15,000 lines) is available at https://github.com/kevinvallier/TSE_Formal.
Remark 26.1 (Verification Architecture).
The formalization follows the logical dependencies of the theory:
@@ B.6 Unification of Discount Factors
This subsection establishes the relationship between fundamental discount rates and derived quantities used throughout the paper. [^2] These results have been formally verified in Lean 4; see Section 26.1.
@@ E.2 Proof of Slack Budget Lemma
E.3 Gronwall’s Inequality
Lemma E.1 (Gronwall).